OpenStudy (anonymous):

5th grade math help!!! Draw a circle.Shae all but 1/6 of it. What percent of the circle is NOT shaded?

OpenStudy (anonymous):

1/6X100

OpenStudy (anonymous):



OpenStudy (anonymous):

now what

OpenStudy (mattfeury):

If you have one whole circle, let's call it's size 1. \[1 = \frac{6}{6}\]If 5/6 of that are shaded, then 1/6 is not shaded.

OpenStudy (anonymous):

?

OpenStudy (mattfeury):

If you have a pizza with 6 slices, and 5 are pepperoni and 1 is sausage then... 5/6 of the pizza is pepperoni, and 1/6 of the pizza is sausage. Does that make sense?

OpenStudy (anonymous):

"What percent of the circle is NOT shaded?" 1/6 not shaded = 0.167 percent? 0.167*100=16.7%

OpenStudy (anonymous):

im in 5th grade member this

OpenStudy (mattfeury):

Do you understand what I said about 6 slices of pizza?

OpenStudy (anonymous):

no

OpenStudy (mattfeury):

Ok, say we have a pizza cut into 6 slices... 

OpenStudy (mattfeury):

If 5 of these slices are pepperoni, then we can say this pizza is 5/6 ("five sixths") pepperoni. 5 pepperoni out of 6 total.

OpenStudy (mattfeury):

Make sense so far?